OpenStudy (anonymous):

A stockbroker earns a commission by a flat fee for each transaction as well as a fee per share. On a particular transaction, the stockbroker sold 5,000 shares to a client at a rate of $0.01 per share. The cost of the transaction totaled $57.50. Which of the following equations describes the cost of this transaction in terms of shares sold?

OpenStudy (anonymous):

\[0.01*5000 = 57.50 -x\] Where x is the transactionfee

OpenStudy (anonymous):

"Which of the following" are there any options given ?

OpenStudy (anonymous):

I missread it's \[t=0.01x+7.5\] t= total x= shares

OpenStudy (anonymous):

A. y = 0.01x B. y = 0.01x + 7.5 C. y = 0.01x + 4,999.425 D. y = 0.01x + 5,000

OpenStudy (anonymous):

It's B then

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

np